trentin:

Solve for x: |2x + 6| − 4 = 20

dude:

First you want to add 4 to both sides

trentin:

24

dude:

So we know that we can have 2 solutions \[2x+6=-24\quad \quad \mathrm{or}\quad \:\quad \:2x+6=24\]

dude:

We just solve for x on both equations; #1 \[2x+6=−24\] Subtract 6 \[2x=-30\] Divide by 2 on both sides x=-15

dude:

#2 \[2x+6=24\] Subtract 6 on both sides \[2x=18\] Divide by 2 on both sides \[x=9\]

trentin:

x=15,9

dude:

-15*

dude:

And those are your solutions \[x=-15, x=9\]

trentin:

yes